# Training a Huggingface Model on AWS Sagemaker (Without Tears)

### v1

@@ 2.2.1 Your First Huggingface Script on AWS Sagemaker
Figure LABEL:working-sample (in Appendix) shows a working script adapted from the Huggingface sample code to launch a Python script that uses the transformers library with the sagemaker.huggingface.HuggingFace model estimator.
While the script is running, users can navigate to the Amazon SageMaker > Training > Training jobs page on the left panel and track the progress of the model training. For example, Figure 3 shows the dashboard of the successful/failed runs of the scripts that were launched using the sagemaker.huggingface.HuggingFace estimators as presented in Figure LABEL:working-sample.
@@
Up till this point, pointers and procedures presented in the above sections were unavailable in a single point of documentation; bits and pieces lurk around open tutorials, demonstrations and online courses. After a user manages to get a script shown in Figure LABEL:working-sample working, there are many other open and accessible demonstrations to use the various Huggingface libraries with AWS Sagemaker. Notably, we refer readers to start with https://huggingface.co/docs/sagemaker/train for further reading on the generic use of Huggingface models with AWS Sagemaker.
@@
For the rest of this demo paper, we will explain the sample code from Figure LABEL:working-sample and walk through the code line-by-line with the goal of empowering users to understand what they mean and how they should modify them to their specific NLP needs. As we explain the script, hopefully, it will answer Frequently Asked Questions (FAQs) on using Huggingface libraries with AWS Sagemaker. We will also highlight lesser-known AWS Sagemaker features to maximize developers’ productivity by emulating a researchers’ process to use a Jupyter Notebook on non-managed bare-metal virtual machines.
@@ 3.1 Specifying Code Location for HuggingFace Estimator
The snippets above with the fully functional code on Figure LABEL:working-sample (in Appendix) is sufficient to run an actual model training. There are more tips and tricks on Controlling AWS Instance for HuggingFace Estimator and Hyperparameter tuning jobs with Sagemaker in Appendix B and C.

### v2

@@ 2.2.1 Your First Huggingface Script on AWS Sagemaker
Figure 4 (in Appendix) shows a working script adapted from the Huggingface sample code to launch a Python script that uses the transformers library with the sagemaker.huggingface.HuggingFace model estimator.
While the script is running, users can navigate to the Amazon SageMaker > Training > Training jobs page on the left panel and track the progress of the model training. For example, Figure 3 shows the dashboard of the successful/failed runs of the scripts that were launched using the sagemaker.huggingface.HuggingFace estimators as presented in Figure 4.
@@
Up till this point, pointers and procedures presented in the above sections were unavailable in a single point of documentation; bits and pieces lurk around open tutorials, demonstrations and online courses. After a user manages to get a script shown in Figure 4 working, there are many other open and accessible demonstrations to use the various Huggingface libraries with AWS Sagemaker. Notably, we refer readers to start with https://huggingface.co/docs/sagemaker/train for further reading on the generic use of Huggingface models with AWS Sagemaker.
@@
For the rest of this demo paper, we will explain the sample code from Figure 4 and walk through the code line-by-line with the goal of empowering users to understand what they mean and how they should modify them to their specific NLP needs. As we explain the script, hopefully, it will answer Frequently Asked Questions (FAQs) on using Huggingface libraries with AWS Sagemaker. We will also highlight lesser-known AWS Sagemaker features to maximize developers’ productivity by emulating a researchers’ process to use a Jupyter Notebook on non-managed bare-metal virtual machines.
@@ 3.1 Specifying Code Location for HuggingFace Estimator
The snippets above with the fully functional code on Figure 4 (in Appendix) is sufficient to run an actual model training. There are more tips and tricks on Controlling AWS Instance for HuggingFace Estimator and Hyperparameter tuning jobs with Sagemaker in Appendix B and C.